Costly Corporation plans a new issue of bonds with a par value of $1000, a maturity of 24 years, and an annual coupon rate of 12.0%. Flotation costs associated with a new debt issue would equal 6.0% of the market value of the bonds. Currently, the appropriate discount rate for bonds of firms similar to Costly is 13.0%. The firm's marginal tax rate is 30%. What will the firm's true cost of debt be for this new bond issue?
Answer #1

Present value of bond is:

=PV(13%,24,12%*1000,1000)

=927.17

Floation cost =6%

Market price after floation cost =927.17*(1-0.06) = 871.541

Bond's cost of debt will be:

=RATE(24,12%*1000,-871.541,1000)

=13.86%

After tax =13.86%(1-0.3) =9.70%